# Formal Semantic Geometry over Transformer-based Variational AutoEncoder

### v2

@@ 7 Conclusion and Future Work
In this study, we investigate the localisation of general semantic features to enhance the controllability and explainability of distributional space from the perspective of formal semantics, which is currently under-explored in the NLP domain. We first propose the formal semantic features as role-content and define the corresponding geometrical framework. Then, we propose a supervision approach to bind the semantic role and word content. In addition, we propose a novel traversal probing approach to assess the latent space geometry based on information set and entropy. We extensively evaluate the latent space geometry through the geometrical operations, such as traversal, arithmetic, and our guided traversal. Experimental results indicate the existence of formal semantic geometry. In the future, we will explore the In-context-learning of explanatory reasoning of LLMs based on our formal semantic geometry framework.
@@ 8 Limitations
1. Limitation of data source: this work only focused on explanatory sentences, such as atomic sentences. Whether the semantic separability of other corpora emerges over latent space is not explored. 2. Role-content clusters overlapping: the geometric analysis indicates that the role-content regions still have significant overlapping, so we can propose a new task, naming “sentence semantic disentanglement”, which is how we can better separate/disentangle the semantic features to provide better localisation or composition behaviour over distributional semantic spaces in Computational Linguistics.
@@ Appendix A Experiment Setting
Table 8 displays the statistical information of the datasets used in the experiment. The data of the two datasets partially overlap, so only the unique explanations are selected as the experimental data. The rationale for choosing explanatory sentences is that they are designed for formal/localised/symbolic semantic inference task in natural language form, which provides a semantically complex and yet controlled experimental setting, containing a both well-scoped and diverse set of target concepts and sentence structures, providing a semantically challenging yet sufficiently well-scoped scenario to evaluate the syntactic and semantic organisation of the space.
@@
Table 9: Cluster Information.

### v3

@@ 7 Conclusion and Future Work
In this study, we investigate the localisation of general semantic features to enhance the controllability and explainability of distributional space from the perspective of formal semantics, which is currently under-explored in the NLP domain. We first propose the formal semantic features as role-content and define the corresponding geometrical framework. Then, we propose a supervision approach to bind the semantic role and word content. In addition, we propose a novel traversal probing approach to assess the latent space geometry based on information set and entropy. We extensively evaluate the latent space geometry through geometrical operations, such as traversal, arithmetic, and our guided traversal. Experimental results indicate the existence of formal semantic geometry.
Since recent theoretical works reveal that the LLMs can encode linear symbolic concepts Jiang et al. (2024b), in the future, we will explore their in-context learning of compositional semantics based on our formal semantic geometry framework.
@@ 8 Limitations
1. Limitation of data source: this work only focused on explanatory sentences. Whether the semantic separability of other corpora emerges over latent space is not explored. 2. Role-content clusters overlapping: the geometric analysis indicates that the role-content regions still have significant overlapping over distributional spaces. Therefore, a new potential task can be how we can better separate/disentangle the semantic features (role-content) to provide better localisation or composition behaviour over distributional semantic spaces in the Computational Linguistics domain, further assisting downstream tasks, such as Natural Language Reasoning, Compositional Generalisation, etc. 3. Large Language Models: this paper only investigates the BERT-GPT2 architecture based on the current state-of-the-art language VAE (Optimus). The larger decoder is out of the scope of this work and needs to be investigated in the future.
Acknowledgements
We appreciate the reviewers for their insightful comments and suggestions. This work was partially funded by the EPSRC grant EP/T026995/1 entitled “EnnCore: End-to-End Conceptual Guarding of Neural Architectures” under Security for all in an AI enabled society, by the Swiss National Science Foundation (SNSF) project NeuMath (200021_204617), by the CRUK National Biomarker Centre, and supported by the Manchester Experimental Cancer Medicine Centre and the NIHR Manchester Biomedical Research Centre.
@@ Appendix A Experiment Setting
Table 8 displays the statistical information of the datasets used in the experiment. The data of the two datasets partially overlap, so only the unique explanations are selected as the experimental data. The rationale for choosing explanatory sentences is that they are designed for formal/localised/symbolic semantic inference task in natural language form, which provides a semantically complex and yet controlled experimental setting, containing a both well-scoped and diverse set of target “concepts” and sentence structures, providing a semantically challenging yet sufficiently well-scoped scenario to evaluate the syntactic and semantic organisation of the space. Besides, those concepts mentioned in the corpus, such as animal is a kind of living thing, are fundamental to human semantic understanding.
@@
Table 9: Cluster Information for explanatory sentences, we use a k-means classifier to classify the sentence representations and manually evaluate each class.